The yield of the reaction is 91.5% and the purity of the oxygen gas is 78.5%.
250.0g of impure oxygen gas react with an excess amount of octane.
Calculate the mass (in grams) of water vapour that will be formed.
Equation: 2C8H18 + 25O2-> 16CO2 + 18H2O(g)
1
Expert's answer
2020-04-23T13:08:58-0400

Find the mass of pure oxygen by the formula:

m2(O2)= (m1(O2)*"\omega"(O2))/100%

m1-impure oxigen

"\omega" - purity of the oxygen


m2(O2)=(250g*78,5%)/100%=195,25g

Find the amount of substance oxygen:

n(O2)=m2(O2)/M(O2) = 196,25g/32g/mol=6,13mol


Then, according to the reaction equation, the amount of substance water will be equal to:


n(H2O)=(6,13mol*18mol)/25mol=4,41mol


Then, theoretical mass of water:

m1(H2O)=n(H2O)*M(H2O)=4,41mol*18g/mol=79,38g


The practical mass of water:

m2(H2O)=(79,38g*91,5%)/100% = 72,63g